OpenStudy (anonymous):

consider a parabola, P, that is congruent to y= x^2 and with vertex (2,-4). find the equation of a new parabola that results if P is translated 3 units to the right and translated 1 unit down.

OpenStudy (amistre64):

.... but, I assumed the vertex was at 0,0

OpenStudy (amistre64):

(y+4) = (x-2)^2 ; now note it lol (y+4+1) = (x-2-3)^2 (y+5) = (x-5)^2